OpenStudy (bloosweater):

Will fan and give medals! Part 1: Using complete sentences, compare the key features and graphs of sine and cosine. What are their similarities and differences? Part 2: Using these similarities and differences, how would you transform f(x) = 2 sin(2x - π) + 3 into a cosine function in the form f(x) = a cos(bx - c) + d? PLEASE HELP ??

OpenStudy (bloosweater):

@Abhisar

OpenStudy (bloosweater):

@perl @triciaal

OpenStudy (triciaal):

@jim_thompson5910 @nikato

OpenStudy (bloosweater):

I already kinda have an answer for the first part, but I need to know if it's right.

OpenStudy (wolf1728):

Go here: http://www.1728.org/trigtutr.htm Scroll almost to the bottom and there you will see a graph of the sine and the cosine function.

OpenStudy (triciaal):

did you include that the max is 1 and the min is -1 follow the same pattern shift pi/4

OpenStudy (bloosweater):

That's the part i already sort of understand, I just need to figure out how to apply it exactly to the second part of the question :L

OpenStudy (bloosweater):

A sine function on a graph is the same as its cosine form, merely moved over π/2 units to the right, because cos (t) = sin (t + π/2). ^that's what i have so far

OpenStudy (bloosweater):

@satellite73 @ganeshie8

OpenStudy (bloosweater):

Figured it out myself. For others with the same question in the future: f(x) = 2 sin(2x - π) + 3 f(x) = 2 cos(2x - π) + 3 f(x) = 2 cos(2x - π - π/2) + 3 f(x) = 2 cos(2x - 3π/2) + 3